find an equation of a line that passes through the point (6,-1) and perpendicular to the line 5x + 2y =-9
Answer by mananth(16946)

5 x + 2 y = -9
Find the slope of this line

2 y = -5 x -9
Divide by 2
y = -2.500 x + -4.5
Compare this equation with y=mx+b
slope m = -2.50 =-5/2

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be 2/5

m= 2/5 ,point ( 6 , -1 )
Find b by plugging the values of m & the point in
y=mx+b
-1 = 2.40 + b
b= -3.4=-17/5
m= 2/5
The required equation is y = 2/ 5 x -17/ 5
m.ananth@hotmail.ca